OpenStudy (anonymous):

F(s)=(s+2)/(s+9) find the max of F(s) on interval [1,4]

OpenStudy (anonymous):

quotient rule first?

OpenStudy (anonymous):

\[F'(s)=\frac{d}{ds}\frac{s+2}{s+9}=\frac{(1)(s+9)-(s+2)(1)}{(s+9)^2}=0 \implies \frac{7}{(s+9)^2}=0\] This is never true for any s. This means there are no local max/min. So test the end points: \[f(1)=\frac{3}{10}, f(4)=\frac{6}{13}\] So 6/13 should be the max. and 3/10 should be the min.

OpenStudy (anonymous):

THANK YOU